OpenStudy (anonymous):

What is a possible first step in eliminating a variable in the following system of equations? x + y = 3 3x – 4y = 23 Choices: A.Multiply the first equation by –4. B.Multiply the first equation by 3. C.Choices 1 and 2. D.None of the above.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

u want to solve only 1 variable

OpenStudy (anonymous):

A. -4x-4y=-12 3x-4y=23 Subtracting the equations eliminates y B. 3x+3y=9 3x-4y=23 Subtracting the equations eliminates x So, choices A and B both seem to work.

OpenStudy (anonymous):

so the answer is C?